OpenStudy (anonymous):

log base 3(1/9) = x What is x?

sam (.sam.):

log^3((1)/(9))=x The logarithm base 3 of (1)/(9) is (1)-log^3(9). x=(1)-log^3(9) The logarithm base 3 of 9 is 2. x=2

Directrix (directrix):

3^x = 1/9 3^x = 3 ^ (-2) x = -2

OpenStudy (anonymous):

One could have just changed the base3 to base10. \[x=\log_{10}(1/9)/\log_{10}3 \] x=-2